# Systematic Constraint Formulation and Collision Free Trajectory Planning Using Space-Time Graphs of Convex Sets

### v1

@@ III-E Space-Time Configuration Constraints
Figure 10: The intersection of all velocity cone constraints remains convex (a). The final control point must lie within this intersection, which is equivalent to just placing a velocity cone around the second-to-last control point (b).
Figure 11: A dynamic obstacle is shown at different times in $\mathbb{R}^{2}$ (a). The initial and final positions of the vertices are put in the space-time configuration space (b). The vertices are connected to form a shape showing its trajectory in $\mathbb{R}^{3}$ (c).
To make this constraint compatible with a convex solver, we can rearrange the constraint as:
$\lVert(\vec{x}_{v,i+1}-\vec{x}_{v,i})\rVert_{xy,2}$ $\leq v_{\text{max}}(\vec{x}_{v,i+1}-\vec{x}_{v,i})_{t}$ (26)
$(\vec{x}_{v,i+1}-\vec{x}_{v,i})_{t}$ $\geq 0$ (27)
Substituting the bilinear variable $\vec{z}_{v,i}$, this becomes:
@@
The following constraint
$\left(\frac{\vec{z}_{v,i+1}}{y_{v}}-\frac{\vec{z}_{v,i}}{y_{v}}\right)_{t}\geq\epsilon$ (30)
can be substituted to ensure that control points maintain a minimum time separation.
@@
III-F Space-Time Obstacles
@@ III-H Optimization Formulation
$y_{v}\in\{0,1\}$ (36c)
$y_{e}\in\{0,1\}$ (36d)
$\sum y_{(a,v)}+\delta_{s}$
$=\sum y_{(v,b)}+\delta_{f}=y_{v}\quad\forall v\in\mathcal{V}$ (36e)
$A_{v}z_{v,i}\leq b_{v}y_{v}.$ (36f)
$z_{v,n}=\sum_{b}\vec{z}_{(v,b),0}\quad\forall v\in\mathcal{V}$ (36g)
$z_{v,n}{-}z_{v,n-1}=\sum_{b}z_{(v,b),\text{diff}}\quad\forall v\in\mathcal{V}$ (36h)
$\lVert(z_{v,i+1}{-}z_{v,i})\rVert_{xy,2}\leq$
$v_{\text{max}}(z_{v,i+1}-z_{v,i})_{t}$ (36i)
$(\vec{z}_{v,i+1}-\vec{z}_{v,i})_{t}\geq y_{v}\epsilon.$ (36j)
This final formulation includes relaxed binary variables $y$ for each vertex and edge, $n$ control point vectors $z_{v,i}$ within each vertex, and additional vectors representing continuity $\vec{z}_{(v,b),0}$ and differentiability $z_{(v,b),\text{diff}}$ for each directed edge. Each of these functions and constraints is convex and is compatible with general convex solvers.
IV Results
@@ IV-A Standard GCS Comparison
Figure 12: The environment shown in 2D (a) and an isometric view of the environment that will be used in the space-time GCS formulation (b).
Figure 13 shows the trajectory generated by the standard GCS formulation. The sets and their associated Bèzier curve control points are color-coded accordingly. The graph of convex sets consists of four sets and eight directed edges. The total computation time was $0.2s$. The optimal distance traveled is $1.0318m$, matching the geometric minimum.
@@
Figure 14 shows the control points generated by the space-time method from a top-down and an isometric view. When viewed from above, the geometric path is identical to the standard GCS solution. The isometric view illustrates the vertical placement of each spline control point over time. The graph of convex sets consists of four sets and eight directed edges. The total computation time was $0.29s$. The optimal distance traveled in 2d is also $1.0318m$, which matches both the standard GCS solution and the true geometric minimum. These results demonstrate that our modified GCS formulation produces trajectories identical to vanilla GCS in static environments, validating the correctness of our implementation.
Figure 14: Resulting sets and trajectory of the space-time GCS shown from the top (a) and in an isometric view (b).
IV-B Dynamic Environment
Figure 15: Resulting sets and collision-free trajectory of the space-time GCS shown from the top (a) and in an isometric view (b).
@@ IV-B Dynamic Environment
Figure 15 shows the control points generated by our method from a top-down and isometric view. When viewed from above, the geometric path moves in a straight line from start to finish. The isometric view shows that the agent moves quickly in the beginning, as evidenced by the trajectory’s slope, to move past the obstacle, then slows down. The graph of convex sets consists of four sets and eight directed edges. The total computation time was $0.52s$. The optimal distance traveled in 2D is also $1.0m$, which is the true geometric minimum distance. This example demonstrates the capability of our approach in generating collision-free, optimal trajectories in the presence of a dynamic obstacle, without relying on an initial guess.
IV-C Cluttered Environment
The previous example shows that our method can find optimal trajectories avoiding a single obstacle. However, in practical scenarios, there are often significantly more obstacles to navigate. This next scenario illustrates how our method performs in a cluttered environment, finding optimal paths through the graphs of convex sets, and presents a heuristic for convex set generation that improves results. We also discuss some of the shortcomings that arise due to the method used to generate the convex sets.
The scenario parameters are as follows. An agent starts at ($0.5m$, $0m$, $0s$) and must move to ($0.5m$, $1m$, $1s$). The agent’s maximum velocity is $3m/s$, allowing it to travel up to three times the distance from start to finish in the allotted $1s$. Twenty obstacles are randomly initialized between ($0m$, $0.2m$) and ($1m$, $0.8m$) with random constant velocities. Figure 16 shows the starting and ending 2D cross sections of this dynamic environment. Figure 17 shows what this looks like in 3D from the side and an isometric view. Finding an initial guess for a trajectory that avoids all the obstacles in this environment would be extremely challenging, if not impossible.
Figure 16: The obstacles move at constant velocity from a random start location to a random end location. Arrows indicate the direction of motion. The start and end of the trajectory are shown in green and red, respectively.
Figure 17: Cluttered environment shown from the top (a) and side view (b).
The free space is decomposed into a graph of convex sets in 3D. The IRIS splitting algorithm generates each convex set by expanding from a seed point to create the largest possible collision-free convex set containing the seed point. The algorithm is seeded by randomly sampling the environment, discarding samples that fall within obstacles or existing sets, and then expanding the graph of convex sets. Figure 18a shows a 3D view of the graph of convex sets after sampling 80 random points and appropriately discarding the samples. In 3D, many of the sets are visually obscured by obstacles or other sets. Figure 18b shows a cross-section of the graph of convex sets at ($t=0.80s$), with sets color-coded and obstacles shown in black. With this random seed, the graph of convex sets consists of fourteen sets and nine directed edges.
Figure 18: 3D view of the graph of convex sets after sampling 80 points (a) and a cross section of the graph at $t=0.80s$ (b).
Figure 19 shows a trajectory generated by our method. There are no collisions with any of the 20 obstacles in the environment. However, while the trajectory is globally optimal with respect to the graph of convex sets, the sets themselves do not provide complete coverage of the free space. As seen in Figure 18b, there is a low density of sets in areas with many obstacles. This results in a trajectory that is not globally optimal with respect to the geometric minimum distance. Figure 19 shows the trajectory in 3D from the top and an isometric view. The total computation time was $4.12s$. The optimal distance traveled is $1.28m$.
Figure 19: A top view of the trajectory generated by our method (a) and a perspective view of the trajectory (b).
We ran the same scenario as above while varying the number of IRIS samples, repeating the experiment 100 times with different seeds. For each run, we recorded the number of sets and edges created in the graph, the final trajectory cost, and the total computation time required to generate the graph and optimize the trajectory. As expected, when the number of samples increased, both the number of sets and edges grew, resulting in a lower average trajectory cost.
Table I summarizes how the average results change with the number of samples used to seed the IRIS algorithm. Computation time increases with the size of the graph, while the distance of the trajectory decreases. The time required to compute the trajectory grows substantially as the graph expands. While some GCS papers propose strategies such as pruning the graph after the first branch-and-bound iteration to reduce computation time, we did not implement these methods in this work. Figure 20 shows an example cross section at $t=0.8s$ after sampling 1000 points to seed the IRIS algorithm, resulting in denser set coverage through the center of the environment and enabling solutions closer to the geometric minimum.
TABLE I: Mean Effect of IRIS Sampling on GCS Trajectory Optimization
Samples Sets Edges Cost (m) Time (s)
80 15.22 22.71 1.24 4.75
100 17.05 26.45 1.23 5.28
250 27.58 52.28 1.12 11.50
500 40.16 95.27 1.05 26.80
1000 55.62 148.44 1.03 82.30
Figure 20: Cross section $t=0.8s$ with 1000 samples used to seed the IRIS algorithm (a). The convex sets are densely packed in the center of the environment, resulting in a shorter trajectory (b).
While the method successfully avoids all obstacles, these results underscore the impact of convex set density on trajectory optimality. Sparse coverage in regions with many obstacles can lead to suboptimal paths, even if the solution is globally optimal within the constructed graph. Increasing the number of samples improves set density and solution quality but comes at the cost of significantly increased computation time.
V Conclusion
@@
We have presented a GCS formulation for trajectory generation in cluttered, dynamic environments. The constraint development strategy introduced here extends the applicability of GCS to a broader range of trajectory optimization problems. Although the convex set generation method limits solution quality, the GCS framework provides safety and optimality guarantees that are rarely available in other trajectory optimization methods. Additionally, the space-time formulation enables GCS to handle dynamic environments effectively. Notably, this approach allows optimal trajectories to be generated without requiring a good initial spatial or temporal guess. Overall, this work demonstrates a strategy to implement GCS constraints, as well as how the GCS framework can be extended to handle dynamic environments.

### v2

@@ 3.5 Space-Time Configuration Constraints
Figure 10: (a) The intersection of all velocity cone constraints remains convex. (b) The final control point must lie within this intersection, which is equivalent to just placing a velocity cone around the second-to-last control point.
Figure 11: (a) A dynamic obstacle is shown at different times in $\mathbb{R}^{2}$. (b) A rendered view of the obstacle in $\mathbb{R}^{3}$.
To make this constraint compatible with a convex solver, we rearrange it to be
$\lVert(\vec{x}_{v,i+1}-\vec{x}_{v,i})\rVert_{xy,2}$ $\leq v_{\text{max}}(\vec{x}_{v,i+1}-\vec{x}_{v,i})_{t},$ (35)
$(\vec{x}_{v,i+1}-\vec{x}_{v,i})_{t}$ $\geq 0.$ (36)
Substituting the bilinear variable $\vec{z}_{v,i}$, this becomes
@@
The constraint can be modified to
$\left(\frac{\vec{z}_{v,i+1}}{y_{v}}-\frac{\vec{z}_{v,i}}{y_{v}}\right)_{t}\geq\epsilon,$ (39)
which ensures that control points maintain a minimum time separation.
@@
3.6 Space-Time Obstacles
@@ 3.8 Optimization Formulation
$y_{v}\in[0,1]$ (45c)
$y_{e}\in[0,1]$ (45d)
$\sum y_{(a,v)}+\delta_{s}=\sum y_{(v,b)}+\delta_{f}=y_{v}$ (45e)
$z_{s,0}=x_{initial}$ (45f)
$z_{f,n}=x_{final}$ (45g)
$A_{v}z_{v,i}\leq d_{v}y_{v}$ (45h)
$\vec{z}_{a,n}=\sum_{b}\vec{p}_{(a,b)}$ (45i)
$\vec{z}_{b,0}=\sum_{a}\vec{p}_{(a,b)}$ (45j)
$\begin{bmatrix}A_{a}\\ A_{b}\end{bmatrix}\vec{p}_{(a,b)}\;\leq\;\begin{bmatrix}\vec{d}_{a}\\ \vec{d}_{b}\end{bmatrix}y_{(a,b)}$ (45k)
$\vec{z}_{a,n-1}=\vec{z}_{a,n}-\sum_{b}\vec{q}_{(a,b)}$ (45l)
$\vec{z}_{b,1}=\vec{z}_{b,0}+\sum_{b}\vec{q}_{(a,b)}$ (45m)
$\left\lVert\vec{q}_{(a,b)}\right\rVert_{2}\leq\gamma y_{(a,b)}$ (45n)
$\lVert(z_{v,i+1}{-}z_{v,i})\rVert_{xy,2}\leq v_{\text{max}}(z_{v,i+1}-z_{v,i})_{t}$ (45o)
$(\vec{z}_{v,i+1}-\vec{z}_{v,i})_{t}\geq y_{v}\epsilon.$ (45p)
This final formulation includes relaxed binary variables $y$ for each vertex and edge, $n$ control point vectors $z_{v,i}$ within each vertex, and additional vectors representing continuity $\vec{p}_{(v,b)}$ and differentiability $\vec{q}_{(v,b)}$ for each directed edge. The variables $z_{s,0}$ and $z_{f,n}$ represent the first control point of the first set and the final control point of the final set, respectively. Each of these functions and constraints is convex and is compatible with general convex solvers.
4 Results
@@ 4.1 Standard GCS Comparison
Figure 12: (a) The environment shown in 2D and (b) an isometric view of the environment that will be used in the space-time GCS formulation.
Figure 13 shows the path generated by the standard GCS formulation. The brightly colored regions are the GCS convex sets. Control points are indicated by blue circles and are connected to adjacent control points by blue lines. The graph of convex sets consists of four sets and eight directed edges. The total computation time was $0.2s$. The optimal distance traveled is $1.0318m$, matching the geometric minimum.
@@
Figure 14 shows the connected Bézier control points generated by the space-time method from a top-down and an isometric view. Control points are indicated by blue circles and are connected to adjacent control points by blue lines. When viewed from above, the geometric path is identical to the standard GCS solution. The isometric view illustrates the vertical placement of each spline control point over time. The graph of convex sets consists of four sets and eight directed edges. The total computation time was $0.29s$. The optimal distance traveled in 2d is also $1.0318m$, which matches both the standard GCS solution and the true geometric minimum. These results demonstrate that our modified GCS formulation produces trajectories identical to vanilla GCS in static environments, validating the correctness of our implementation.
Figure 14: (a) Resulting sets and trajectory of the space-time GCS shown from the top and (b) an isometric view.
4.2 Dynamic Environment
Figure 15: (a) Resulting collision-free trajectory of the space-time GCS shown from the top and (b) in a perspective view. The convex sets are omitted to visualize the trajectory clearly.
@@ 4.2 Dynamic Environment
Figure 15 shows the connected Bézier control points generated by our method from a top-down and perspective view. Control points are indicated by blue circles and are connected to adjacent control points by blue lines. When viewed from above, the geometric path moves in a straight line from start to finish. The perspective view shows that the agent moves quickly in the beginning, as evidenced by the trajectory’s slope, to move past the obstacle, then slows down. The graph of convex sets consists of four sets and eight directed edges. The total computation time was $0.52s$. The optimal distance traveled in 2D is also $1.0m$, which is the true geometric minimum distance. This example demonstrates the capability of our approach in generating collision-free, optimal trajectories in the presence of a dynamic obstacle, without relying on an initial guess.
4.3 Monte-Carlo Simulations
Figure 16: Illustration of the Monte-Carlo simulation scenario with 10 obstacles. The trajectory moves from the start point to the end while avoiding the obstacles moving across the middle.
Figure 17: Rendered result of a minimum distance trajectory moving through a cluttered environment of 14 obstacles. Recall that the vertical direction represents time, meaning that this trajectory avoids spatial and temporal conflicts. This trajectory has a length of $1m$ in the xyx-y plane.
Figure 18: Monte-Carlo trajectory cost distributions.
Figure 19: Percentage of failed trials as a function of obstacle count. Failures peak just below $2.5\%$ out of 1,000 trials.
Figure 20: Count of Monte-Carlo runs achieving a cost within $0.01\%$ of the geometric minimum, as a function of obstacle count. The decrease reflects IRIS coverage limitations and obstacle interference.
To further evaluate the robustness of our space-time GCS formulation, we present a series of Monte-Carlo simulations in cluttered dynamic environments. An agent with state $(x,y,t)$ starts at ($0.5m$, $0m$, $0s$) and must move to ($0.5m$, $1m$, $1s$). The maximum allowable velocity is $3~\text{m/s}$, allowing the agent to traverse up to three times the straight-line distance within the available time.
Obstacles are incrementally introduced, ranging from one to fifteen in total. Each obstacle moves with a piecewise-constant velocity, beginning at evenly spaced positions along the lines $x=0$ and $x=1$, and ending at randomly assigned corresponding destinations on the opposite lines, $x=1$ and $x=0$, respectively. Randomly selecting a corresponding final location forces many of the obstacles to cross in front of each other, resulting in a complicated collision-free space. Obstacles are arranged such that they never collide with each other during their motion. Figure 16 depicts an illustration of the scenario with 10 obstacles. We varied the number of obstacles from 1 to 15 and conducted 1,000 randomized trials for each case. Randomization affects both the obstacle end-locations and the seed points used for IRIS-based convex set generation. To balance computation time and free-space coverage in the graph, we restrict the IRIS algorithm to generate at most $25$ convex sets per trial.
Figure 17 shows a trajectory generated using our formulation moving through a cluttered environment with 14 moving obstacles. Despite the density of obstacles, the agent finds a feasible path in the majority of trials, highlighting the flexibility of the ST-GCS framework.
To quantify performance, we analyze the trajectory cost across all Monte-Carlo runs. Figure 18 presents a box-and-whisker plot of the cost distribution as a function of obstacle count. The horizontal bars represent the median value, the grey boxes represent the inter-quartile range, the vertical black lines represent results above the 25th percentile and below the 75th percentile, and the circles represent outliers. As expected, increasing the number of obstacles raises the median cost, reflecting the reduced availability of direct free-space corridors. We consider trajectories shorter than $1.0001m$, within a percent of a percent of the best possible minimum, to be globally optimal. The median value of each experiment is globally optimal up to and including $11$ obstacles.
We also measure the failure rate of each experiment, defined as the percentage of runs where no feasible trajectory is returned. Failures occur due to a combination of: (1) insufficient connectivity of the graph formed by IRIS, (2) infeasibility arising from there not being a path through the dynamic obstacles, and (3) ill-conditioning in the convex solver. Figure 19 shows that for most scenarios the failure rate was between $1\%$ and $2\%$. Because there is no trend in the error rate with respect to the number of obstacles, we infer that these failures were caused by ill-conditioning in the convex solver. However, because we only measured failure for up to 15 obstacles, this does not mean the error rate may follow a trend as more obstacles are introduced. There were no experiments with a failure rate above $2.5\%$.
Finally, we compare achieved trajectories against the theoretical geometric minimum path length. Figure 20 shows the percentage of trials that we consider to be globally optimal (within $0.01\%$ of the $1m$). This percentage decreases as the number of obstacles increases, likely due to three factors: limited IRIS set coverage, obstruction of direct paths by obstacles, and the finite maximum velocity constraint.
Overall, the Monte-Carlo simulations demonstrate that the proposed ST-GCS framework reliably generates collision-free trajectories in cluttered environments, with failure rates remaining low even under significant obstacle density. While trajectory optimality degrades with insufficient IRIS coverage and increasing clutter, these results confirm the framework’s ability to be applied in dynamic environments.
5 Conclusion
@@
We have presented a GCS formulation for trajectory generation in cluttered, dynamic environments. The constraint development strategy introduced here extends the applicability of GCS to a broader range of trajectory optimization problems. The space-time formulation enables GCS to handle cluttered dynamic environments effectively. Notably, this approach allows optimal trajectories to be generated without requiring a good initial spatial or temporal guess.
Future work includes accounting for uncertainty in the positions of obstacles as well as adapting the space-time formulation to account for real vehicle dynamics. Overall, this work demonstrates a strategy to implement GCS constraints, as well as how the GCS framework can be extended to handle dynamic environments.
@@ Acknowledgment
Declaration of generative AI and AI-assisted technologies in the writing process.
During the preparation of this work, the authors used ChatGPT to improve the readability and wording of different sections of the paper. After using ChatGPT, the authors extensively reviewed and edited the content to ensure accuracy and coherence. We take full responsibility for the content of the published article.
[…]